# Free Doubly-Infinitary Distributive Categories are Cartesian Closed

### v6

@@ 4.1. The $\mathbf{Fam}$ example
We recall that $\mathbf{Fam}(\mathbb{C})$ is extensive for any category $\mathbb{C}$. However, $\mathbf{Fam}(\mathbb{C})$ can lack products and, hence, it might not be doubly-infinitary distributive We, however, have:
Example 4.1.
$\mathbf{Fam}{\mathbb{C}}$ is always doubly-infinitary distributive provided that the category $\mathbb{C}$ has products. Indeed, $\mathbf{Fam}$ lifts to a pseudomonad on $\mathbf{ProdCat}$, whose pseudoalgebras are the doubly-infinitary distributive categories. $\mathbf{Fam}{\mathbb{C}}$ is precisely the free $\mathbf{Fam}$-pseudoalgebra ovn the product-complete category $\mathbb{C}$.
4.2. The category of sets
@@
Example 4.3.
Similarly, any category of presheaves $[\mathbb{D}^{op},\mathbf{Set}]$ is doubly-infinitary distributive, seeing that both products and coproducts are computed pointwise (i.e., the evaluation functors create products and coproducts). More generally, the category of functors $[\mathbb{D}^{op},\mathbb{C}]$ is doubly-infinitary distributive provided that $\mathbb{C}$ is so.
Example 4.4.
The category $\mathbf{Fam}\left(\mathbf{Set}^{\mathrm{op}}\right)=\mathbf{Dist}(\mathbb{1})$ is a free doubly-infinitary distributive category. This category is commonly studied under the name of the category of polynomials or containers [1]. In particular, Theorem 2.4 implies that the category of polynomials is cartesian closed, as was previously noted in [5].
4.3. $\mathbf{ProdCat}$ is bicategorically semi-additive
In order to proceed with examples, we remark that the $2$-category $\mathbf{ProdCat}$ has a natural enrichment over the $2$-category of symmetric monoidal categories with the multilinear multicategorical structure, and has products. This shows that $\mathbf{ProdCat}$ is bicategorically semi-additive (see, for instance, [25] for the notion of semi-additive category).
Lemma 4.5.
The $2$-category $\mathbf{ProdCat}$ has a natural enrichment over the $2$-category of symmetric monoidal categories with the multilinear multicategorical structure. Moreover, $\mathbf{ProdCat}$ has bicategorical products.
Proof.
$\mathbf{ProdCat}\to\mathbf{Cat}$
is pseudomonadic and, hence, creates bicategorical products.
The pointwise product of functors gives the monoidal structure on the hom-categories of $\mathbf{ProdCat}$. ∎
As a consequence, $\mathbf{ProdCat}$ has bicategorical biproducts, meaning that bicategorical products and coproducts coincide up to equivalence (see, for instance, [44, 45] and [30, 3.8] for bilimits). In other words:
Lemma 4.6.
$\mathbf{ProdCat}$ has bicategorical coproducts. Moreover, the bicategorical coproducts are given by the bicateogorical products in $\mathbf{ProdCat}$.
With this observation, we can proceed and construct new examples of doubly-infinitary distributive categories. In particular, we have:
Example 4.7.
The free doubly-infinitary distributive category on two objects admits the following description:
$\mathbf{Dist}\left(\mathbb{1}\sqcup_{\mathbf{Cat}}\mathbb{1}\right)$ $\cong\mathbf{Fam}\left(\mathbf{Fam}\left(\left(\mathbb{1}\sqcup_{\mathbf{Cat}}\mathbb{1}\right)^{op}\right)^{op}\right)$
$\cong\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\sqcup_{\mathbf{ProdCat}}\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\right)$
$\cong\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\times\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\right)$
$\cong\mathbf{Fam}\left(\mathbf{Set}^{op}\times\mathbf{Set}^{op}\right),$
where we use the facts that:
  1. (1)
    the functor $op\circ\mathbf{Fam}\circ op:\mathbf{Cat}\to\mathbf{ProdCat}$ is a left biadjoint and thus preserves $2$-categorical coproducts, and
  2. (2)
    bicategorical coproducts and products coincide in $\mathbf{ProdCat}$.
@@ 4.4. Posets
There is an extensive literature on completely distributive lattice, e.g. [17]. We observe that a poset $\mathbb{C}$ is a doubly-infinitary distributive category if and only if it is a completely distributive lattice.
If a poset is extensive, it is, in particular, a distributive lattice and $a<b$ implies that $a=a\wedge b=\bot$, due to disjointness of coproducts in an extensive category. Hence only the trivial poset is extensive.
This shows that doubly-infinitary distributive categories may fail to be extensive.
@@ 4.5. Categories of categorical structures
Consider the category $\mathbf{Cat}$ of small categories and functors. Observe that $\mathbf{Cat}\simeq\mathbf{Fam}{\mathbf{ConCat}}$ for the full subcategory $\mathbf{ConCat}$ of connected categories (i.e., categories $\mathbb{C}$ such that the free groupoid on $\mathbb{C}$ is connected; see, for instance, [28, 1.2]). Seeing that $\mathbf{ConCat}$ has products, $\mathbf{Cat}$ is doubly-infinitary distributive, by Example 4.1. The same argument works to show that the category $\mathbf{Pos}$ of posets and $\mathbf{\boldsymbol{\omega}CPO}$ of $\omega$-chain cocomplete partial orders and $\omega$-cocontinuous functors are doubly infinitary distributive.
@@ 4.6. Category of topological spaces
It is well-known that the category $\mathbf{Top}$ of topological spaces and continuous functions is infinitary distributive in the usual sense: finite products distributive over (potentially infinite) coproducts. However, infinite products in $\mathbf{Top}$ are not quite as well-behaved, making the doubly-infinitary distributive law fail, in general.
Indeed, to study the continuity of the distributor $d$, observe the following. Let $U$ be an open subset of $\coprod_{f:\Pi_{j\in J}I_{j}}\prod_{j\in J}C_{f_{j}j}$. Then, by definition of the coproduct topology, $U=\coprod_{f\in\Pi_{j\in J}I_{j}}U_{f}$ for open subsets $U_{f}\subseteq\prod_{j\in J}C_{f(j)j}$. By definition of the product topology, $U_{f}=\bigcup_{\alpha\in A_{f}}\bigcap_{\beta\in B_{f\alpha}}\pi^{-1}_{j_{f\alpha\beta}}(V_{f\alpha\beta})$, where $V_{f\alpha\beta}$ is an open subset of $C_{f(j_{f\alpha\beta})j_{f\alpha\beta}}$, $A_{f}$ is some potentially infinite set and $B_{f\alpha}$ is a finite set. Now,
$d^{-1}(U)$ $=d^{-1}(\coprod_{f\in\Pi_{j\in J}I_{j}}U_{f})$
$=d^{-1}(\coprod_{f:\Pi_{j\in J}I_{j}}\bigcup_{\alpha\in A_{f}}\bigcap_{\beta\in B_{f\alpha}}\pi^{-1}_{j_{f\alpha\beta}}(V_{f\alpha\beta}))$
$=[\langle\iota_{f(j)}\circ\pi_{j}\mid j\in J\rangle\mid f\in\Pi_{j\in J}I_{j}](\coprod_{f\in\Pi_{j\in J}I_{j}}\bigcup_{\alpha\in A_{f}}\bigcap_{\beta\in B_{f\alpha}}\pi^{-1}_{j_{f\alpha\beta}}(V_{f\alpha\beta}))$
$=\bigcup_{f:\Pi_{j\in J}I_{j}}\langle\iota_{f(j)}\circ\pi_{j}\mid j\in J\rangle(\bigcup_{\alpha\in A_{f}}\bigcap_{\beta\in B_{f\alpha}}\pi^{-1}_{j_{f\alpha\beta}}(V_{f\alpha\beta}))$
$=\bigcup_{f:\Pi_{j\in J}I_{j}}\bigcup_{\alpha\in A_{f}}\langle\iota_{f(j)}\circ\pi_{j}\mid j\in J\rangle(\bigcap_{\beta\in B_{f\alpha}}\pi^{-1}_{j_{f\alpha\beta}}(V_{f\alpha\beta}))$
@@
$=\bigcup_{f\in\Pi_{j\in J}I_{j}}\bigcup_{\alpha\in A_{f}}\bigcap_{\beta\in B_{f\alpha}}\langle\iota_{f(j)}\circ\pi_{j}\mid j\in J\rangle(\prod_{j\in J}\left\{\begin{array}[]{ll}V_{f\alpha\beta}&\text{if $j=j_{f\alpha\beta}$}\\ C_{f(j_{f\alpha\beta})j_{f\alpha\beta}}&\text{otherwise}\end{array}\right\})$
$=\bigcup_{f:\Pi_{j\in J}I_{j}}\bigcup_{\alpha\in A_{f}}\bigcap_{\beta\in B_{f\alpha}}\prod_{j\in J}\left\{\begin{array}[]{ll}\iota_{f(j_{f\alpha\beta})}(V_{f\alpha\beta})&\text{if $j=j_{f\alpha\beta}$}\\ \iota_{f(j)}(C_{f(j)j})&\text{otherwise}\end{array}\right.$
Meanwhile, the open sets of $\prod_{j\in J}\coprod_{i\in I_{j}}C_{ij}$, by definition of the product and coproduct topologies, are the sets of the form
$\bigcup_{k\in K}\bigcap_{l\in L}\pi^{-1}_{j_{kl}}\iota_{i_{kl}}(W_{kl})=\bigcup_{k\in K}\bigcap_{l\in L}\prod_{j\in J}\left\{\begin{array}[]{ll}\iota_{i_{kl}}(W_{kl})&\text{if $j=j_{kl}$}\\ \coprod_{i\in I_{j}}C_{ij}&\text{otherwise}\end{array}\right.$
for some set $K$, finite set $L$ and open subsets $W_{kl}$ of $C_{i_{kl}j_{kl}}$. This shows that the map $d$ above is continuous if $J$ is finite. However, if $J$ is infinite, taking $I_{j}=\{0,1\}$, $C_{ij}=\mathbb{R}$, $A_{\lambda\_.0}=\{*\}$ and $A_{f}=\emptyset$ for all other $f$, $B_{f\alpha}=\{*\}$, and $V_{f**}=\mathbb{R}$ gives a counter-example to continuity. Indeed, then $U=\iota_{\lambda\_.0}(\coprod_{j\in J}\mathbb{R})$ is an open subset of $\prod_{f\in J\to\{0,1\}}\coprod_{j\in J}\mathbb{R}$ and $d^{-1}(U)=\coprod_{j\in J}\iota_{0}(\mathbb{R})\subseteq\coprod_{j\in J}\prod_{i\in\{0,1\}}\mathbb{R}$ is not open.
We see that $\mathbf{Top}$ is infinitary distributive (actually lextensive), but not doubly-infinitary distributive.
Remark 4.8 (Cantor space).
The proof that the category of topological spaces is not doubly-infinitary distributive can be abridged by the fact that the Cantor space $\mathbb{K}$ is an infinite compact topological space, and, hence, not discrete.
Specifically, $\mathbb{K}$ can be constructed as the infinite product $\prod_{j\in\mathbb{N}}\coprod_{i\in\left\{0,1\right\}}1$ in the category $\mathbf{Top}$ of topological spaces, where $1$ represents the space with only one point. Since $\mathbb{K}$ is compact by Tychonoff’s theorem, it cannot be homeomorphic to an infinite coproduct of non-trivial spaces. More explicitly, the canonical morphism 3.1, in this case, is the morphism $\underline{\mathbb{K}}\to\mathbb{K}$ between the underlying discrete topological space $\underline{\mathbb{K}}$ and $\mathbb{K}$, which is evidently not invertible.
@@
Analogously, we can conclude that, despite being (l)extensive, the category of locales $\mathbb{Loc}=\mathbb{Frm}^{\mathrm{op}}$ is not doubly-infinitary distributive (see [40, Chapter 13] for connectedness in pointfree topology).
4.7. Cartesian closedness vs. doubly-infinitary distributivity
Herein, we establish the comparison between doubly-infinitary distributivity and cartesian closedness.
@@
Let $\mathbb{C}$ be infinitary distributive category. The category $\mathbb{C}$ is cartesian closed if and only if $\mathbf{Fam}(\mathbb{C})$ is cartesian closed.
@@
It is well-known that, if $\mathbb{C}$ is cartesian closed, so is $\mathbf{Fam}(\mathbb{C})$ (we refer the reader to [3, 2] for instance).
Reciprocally, if $\mathbf{Fam}(\mathbb{C})$ is cartesian closed, we know that the canonical inclusion
$\mathbb{C}\to\mathbf{Fam}(\mathbb{C})$
is fully faithful, and has the left adjoint given by the coproduct functor $\coprod$ (see (3.2)). By the definition of infinitary distributive categories, we also know that $\coprod$ preserves finite products. Then, by [22, Proposition 4.3.1], $\mathbb{C}$ forms an exponential ideal in $\mathbf{Fam}(\mathbb{C})$. In particular, $\mathbb{C}$ is cartesian closed. ∎
This shows that, whenever $\mathbb{C}$ is an example of (doubly-)infinitary distributive category that is not cartesian closed, $\mathbf{Fam}(\mathbb{C})$ is another such an example.
With the observation above, we can see that doubly-infinitary distributive categories are not necessarily cartesian closed. More precisely, given an infinitary distributive category $\mathbb{C}$ that is not cartesian closed, $\mathbf{Fam}(\mathbb{C})$ is not cartesian closed and, yet, it is doubly-infinitary distributive. For instance:
@@
$\mathbf{Fam}(\mathbf{Top})$ is doubly-infinitary distributive, as a free coproduct completion of a product-complete category.
At the same time, $\mathbf{Top}$ is well-known not to be cartesian closed, as the product $\mathbb{S}\times-$ with the Sierpinski space $\mathbb{S}$ does not preserve the coequalizer of some parallel pair $f,g:X\to Y$ [7, Proposition 7.1.2].
Since $\mathbf{Top}$ is infinitary distributive, we conclude that $\mathbf{Fam}(\mathbf{Top})$ is not cartesian closed by Theorem 4.10.
@@
Consider the category $\mathbf{ConTop}$ of locally connected connected topological spaces, and continuous functions. Observe $\mathbf{Fam}(\mathbf{ConTop})$ is, up to equivalence, the category $\mathbf{Fam}(\mathbf{ConTop})\simeq\mathbf{LocConTop}$ of locally connected topological spaces and continuous functions (see, for instance, [8, Proposition 6.15]). $\mathbf{ConTop}$ has products, given by the usual product topology. Therefore, $\mathbf{LocConTop}$ is doubly-infinitary distributive, as an instance of Example 4.1. Again, $\mathbf{LocConTop}$ is not cartesian closed, by the same argument as for $\mathbf{Top}$.
We can generalize Example 4.12 and the examples of 4.5 by making use of the notion of connected objects (see, for instance, [8, Definition 6.1.3]). An object $C$ in a category $\mathbb{C}$ is connected if the hom-functor $\mathbb{C}\left(C,-\right)$ preserves coproducts. We, then, define the full subcategory $\mathsf{Con}\mathbb{C}$ of the connected objects in $\mathbb{C}$. We get that:
@@
By [8, Proposition 6.15], we conclude that $\mathbb{C}$ is equivalent to $\mathbf{Fam}(\mathsf{Con}\mathbb{C})$. By Example 4.1, we conclude that $\mathbb{C}$ is doubly-infinitary distributive, since $\mathsf{Con}\mathbb{C}$ has products. ∎
@@
Recall that the usual category $\mathbf{\boldsymbol{\omega}CPO}$ of $\omega$-cpos is not locally cartesian closed (see, for example, [4, Proposition 4]). Therefore, there is some $\omega$-cpo $X$ (concretely, we can take $X$ to be the extended natural numbers $\mathbb{N}_{\infty}$ with the linear order $\leq$), such that the slice category $\mathbf{\boldsymbol{\omega}CPO}/X$ is not cartesian closed. For example, the internal hom $((\mathbb{N},=)\hookrightarrow(\mathbb{N}_{\infty},\leq))\Rightarrow(-)$ does not exist, for the inclusion of the natural numbers with the discrete order into the extended natural numbers with the linear order. Meanwhile, products in $\mathbf{\boldsymbol{\omega}CPO}/X$ are simply given by wide pullbacks in $\mathbf{\boldsymbol{\omega}CPO}$, i.e. the wide pullback in $\mathbf{Set}$ over $|X|$ equipped with the product order. Coproducts in $\mathbf{\boldsymbol{\omega}CPO}/X$ are simply given by coproducts in $\mathbf{\boldsymbol{\omega}CPO}$, i.e. disjoint unions of sets and orders, with the required morphisms to $X$ induced by the cotupling. Therefore, the isomorphism
$\prod_{j\in J}\coprod_{i\in I_{j}}Y_{ij}\cong\coprod_{f\in\prod_{j\in J}I_{j}}\prod_{j\in J}Y_{f_{j}j}$
restricts, if we write $\prod^{X}$ and $\coprod^{X}$ for the product and coproduct in $\mathbf{\boldsymbol{\omega}CPO}/X$, to one
$\prod_{j\in J}{}^{X}\coprod_{i\in I_{j}}{}^{X}(y_{ij}:Y_{ij}\to X)\cong\coprod_{f\in\prod_{j\in J}I_{j}}{}^{X}\prod_{j\in J}{}^{X}(y_{f(j)j}:Y_{f_{j}j}\to X),$
showing that $\mathbf{\boldsymbol{\omega}CPO}/X$ is doubly-infinitary distributive.
Cartesian closed categories are not necessarily doubly-infinitary distributive
Since the functor $A\times-$ is left adjoint in cartesian closed categories, it is well-known that cartesian closed categories are infinitary distributive.

### v7

@@ 5.1. The $\mathbf{Fam}$ example
We recall that $\mathbf{Fam}(\mathbb{C})$ is extensive for any category $\mathbb{C}$. However, $\mathbf{Fam}(\mathbb{C})$ can lack products and, hence, it might not be doubly-infinitary distributive. We, however, have:
Lemma 5.1.
$\mathbf{Fam}{\left(\mathbb{C}\right)}$ is doubly-infinitary distributive provided that the category $\mathbb{C}$ has products.
Proof.
Indeed, $\mathbf{Fam}$ lifts to a pseudomonad on $\mathbf{ProdCat}$, whose pseudoalgebras are the doubly-infinitary distributive categories.
That is to say, $\mathbf{Fam}{\left(\mathbb{C}\right)}$ is precisely the free doubly-infinitary distributive category on the product-complete category $\mathbb{C}$. ∎
5.2. The category of sets and beyond
@@
Example 5.3 (Presheaves and Functor Categories).
Similarly, any category of presheaves $\mathbf{Cat}[\mathbb{D}^{op},\mathbf{Set}]$ is doubly-infinitary distributive, seeing that both products and coproducts are computed pointwise (i.e., the evaluation functors create products and coproducts). More generally, the category of functors $\mathbf{Cat}[\mathbb{D}^{op},\mathbb{C}]$ is doubly-infinitary distributive provided that $\mathbb{C}$ is so.
Example 5.4 (Category of Polynomials).
The category $\mathbf{Fam}\left(\mathbf{Set}^{op}\right)=\mathbf{Dist}(\mathbb{1})$ is a free doubly-infinitary distributive category. This category is commonly studied under the name of the category of polynomials or containers [1]. In particular, Theorem 2.4 implies that the category of polynomials is cartesian closed, as was previously noted in [5].
@@ 5.3. Posets
We observe that a poset $\mathbb{C}$ is a doubly-infinitary distributive category if and only if it is a completely distributive lattice, e.g. [19] for completely distributive lattices.
Moreover, only the trivial poset is extensive. Indeed, if a poset $X$ is extensive, the coproducts of $X$ are disjoint. Hence $a=a\wedge a=\bot$ for any $a\in X$.
This shows that any non-trivial completely distributive lattice provides an example of a category that is not extensive, yet is doubly-infinitary distributive.
@@ 5.4. Categories of categorical structures
Consider the category $\mathbf{cat}$ of small categories and functors. We observe that $\mathbf{cat}\simeq\mathbf{Fam}{\left(\mathbf{Concat}\right)}$ where we denote by $\mathbf{Concat}$ the full subcategory of connected categories in $\mathbf{cat}$ (i.e., small categories $\mathbb{C}$ such that the free groupoid on $\mathbb{C}$ is connected; e.g. [31, 1.2]).
Since $\mathbf{Concat}$ has products, $\mathbf{cat}\simeq\mathbf{Fam}{\left(\mathbf{Concat}\right)}$ is doubly-infinitary distributive by Lemma 5.1. The same argument works to show that the category $\mathbf{Pos}$ of posets and $\mathbf{\boldsymbol{\omega}CPO}$ of $\omega$-chain cocomplete partial orders and $\omega$-cocontinuous functors are doubly-infinitary distributive.
@@ 5.5. Category of topological spaces
It is well known that the category $\mathbf{Top}$ of topological spaces and continuous functions is infinitary distributive in the usual sense: finite products distribute over (possibly infinite) coproducts. However, as we show below, $\mathbf{Top}$ is not doubly-infinitary distributive. This can be seen by considering the so-called Cantor spaces.
More precisely, we examine the instance of the canonical comparison (3.1) with $J=\mathbb{N}$, $I_{j}=2$ for all $j\in\mathbb{N}$, and $C_{ji}=1$ (the one-point space) for all $i\in 2$, $j\in\mathbb{N}$. This yields the continuous map
@@
Since $\prod\limits_{j\in\mathbb{N}}1\cong 1$, the domain of $d$ is the topological sum $\coprod\limits_{r\in 2^{\mathbb{N}}}1$, that is, the discrete space on the underlying set $2^{\mathbb{N}}$. The discrete space $2^{\mathbb{N}}$ is clearly not compact as it is an infinite discrete space.
The codomain of $d$ is the countable product $\prod\limits_{j\in\mathbb{N}}2$ of the discrete two-point space, that is, the Cantor space $\mathbb{K}\cong 2^{\mathbb{N}}$ with the product topology, which is compact by Tychonoff’s theorem, e.g. [11]. The underlying function of $d$ is the identity on $2^{\mathbb{N}}$, but $d$ is not invertible in the category of topological spaces since it fails to reflect compactness.
Hence, $\mathbf{Top}$ is infinitary distributive (indeed lextensive), but not doubly-infinitary distributive.
@@
Analogously, we can conclude that, despite being (l)extensive, the category of locales $\mathbb{Loc}=\mathbb{Frm}^{\mathrm{op}}$ is not doubly-infinitary distributive (see [45, Chapter 13] for connectedness in pointfree topology).
5.6. Doubly-infinitary distributive categories that are not cartesian closed
Through several examples, we establish below that there exist doubly-infinitary distributive categories which are not cartesian closed. This demonstrates that, although freely generated doubly-infinitary distributive categories are cartesian closed, doubly-infinitary distributivity does not imply cartesian closedness.
We begin by examining the relationship between the cartesian closedness of a category $\mathbb{C}$ and that of its free coproduct completion. More precisely:
@@ 5.6. Doubly-infinitary distributive categories that are not cartesian closed
Let $\mathbb{C}$ be an infinitary distributive category. Then $\mathbb{C}$ is cartesian closed if and only if $\mathbf{Fam}(\mathbb{C})$ is cartesian closed.
@@
It is well known that if $\mathbb{C}$ is cartesian closed, then so is $\mathbf{Fam}(\mathbb{C})$ (see, for instance, [3, 2]).
Conversely, suppose that $\mathbf{Fam}(\mathbb{C})$ is cartesian closed. The canonical inclusion
$\mathbb{C}\longrightarrow\mathbf{Fam}(\mathbb{C})$
is fully faithful and admits a left adjoint given by the coproduct functor $\coprod$ (see (3.2)). By the definition of infinitary distributive categories, $\coprod$ preserves finite products. Then, by [25, Proposition 4.3.1], it follows that $\mathbb{C}$ forms an exponential ideal in $\mathbf{Fam}(\mathbb{C})$. Hence, $\mathbb{C}$ is cartesian closed. ∎
This shows that, whenever $\mathbb{C}$ is an example of (doubly-)infinitary distributive category that is not cartesian closed, $\mathbf{Fam}(\mathbb{C})$ is another such an example. More precisely:
Corollary 5.7.
Let $\mathbb{C}$ be a product-complete, infinitary distributive category which is not cartesian closed. Then:
  1. (1)
    $\mathbf{Fam}(\mathbb{C})$ is doubly-infinitary distributive; and
  2. (2)
    $\mathbf{Fam}(\mathbb{C})$ is not cartesian closed.
Proof.
Since $\mathbb{C}$ is product-complete, then $\mathbf{Fam}(\mathbb{C})$ is doubly-infinitary distributive.
Since $\mathbb{C}$ is infinitary distributive and not cartesian closed, $\mathbf{Fam}(\mathbb{C})$ is not cartesian closed. ∎
There are several examples of infinitary distributive categories that are not cartesian closed. For instance:
@@
It is well known that $\mathbf{Top}$ is not cartesian closed, since the product $\mathbb{S}\times$- with the Sierpiński space $\mathbb{S}$ fails to preserve the coequalizer of certain parallel maps $f,g:X\to Y$ [7, Proposition 7.1.2].
Therefore, as $\mathbf{Top}$ is infinitary distributive, we conclude from Corollary 5.7 that $\mathbf{Fam}(\mathbf{Top})$ is not cartesian closed.
@@
Consider the category $\mathbf{ConTop}$ of locally connected topological spaces, and continuous functions. Observe $\mathbf{Fam}(\mathbf{ConTop})$ is, up to equivalence, the category $\mathbf{Fam}(\mathbf{ConTop})\simeq\mathbf{LocConTop}$ of locally connected topological spaces and continuous functions (see, for instance, [8, Proposition 6.15]). $\mathbf{ConTop}$ has products, given by the usual product topology. Therefore, $\mathbf{LocConTop}$ is doubly-infinitary distributive by Lemma 5.1.
Despite being doubly-infinitary distributive, $\mathbf{LocConTop}$ is not cartesian closed.
We can generalize Example 5.9 and the examples of 5.4 by making use of the notion of connected objects (see, for instance, [8, Definition 6.1.3]). An object $C$ in a category $\mathbb{C}$ is connected if the hom-functor $\mathbb{C}\left(C,-\right)$ preserves coproducts. We, then, define the full subcategory $\mathsf{Con}\mathbb{C}$ of the connected objects in $\mathbb{C}$. We get that:
@@
By [8, Proposition 6.15], we conclude that $\mathbb{C}$ is equivalent to $\mathbf{Fam}(\mathsf{Con}\mathbb{C})$. By Lemma 5.1, we conclude that $\mathbb{C}$ is doubly-infinitary distributive, since $\mathsf{Con}\mathbb{C}$ has products. ∎
@@
Recall that the usual category $\mathbf{\boldsymbol{\omega}CPO}$ of $\omega$-cpos is not locally cartesian closed (see, for example, [4, Proposition 4]). For instance, by assuming that the extended natural numbers $\mathbb{N}_{\infty}$ with the linear order, $\mathbf{\boldsymbol{\omega}CPO}/\mathbb{N}_{\infty}$ is not cartesian closed.
Meanwhile, for any object $X$ of $\mathbf{\boldsymbol{\omega}CPO}$, we can conclude that $\mathbf{\boldsymbol{\omega}CPO}/X$ is doubly-infinitary distributive by Lemma 5.10. This shows that $\mathbf{\boldsymbol{\omega}CPO}/\mathbb{N}_{\infty}$ is an example of doubly-infinitary distributive category that is not cartesian closed.
5.7. Cartesian closed categories are not necessarily doubly-infinitary distributive
Since the functor $A\times$- is left adjoint in cartesian closed categories, it is well-known that any coproduct-complete cartesian closed categories are infinitary distributive.